OpenStudy (anonymous):

Determine the area of the region between the given curves: f(x) = x^3 + 3, x = -2 and y = 11 .....show work please

OpenStudy (anonymous):

do you know what is integral?

OpenStudy (anonymous):

That is all the info given.

OpenStudy (anonymous):

well if he don't know what's integral...

OpenStudy (anonymous):

lets assume its f(x)

myininaya (myininaya):

\[\int\limits_{L }^{U} (t op function-bottom function) dx\]

myininaya (myininaya):

top function is y=11 bottom function is x^3+3

OpenStudy (anonymous):

\[\text{area}=\int_{-1}^a \left[11-f(x)\right]dx\]where a is the x value of the point of intersection between f(x) and your horizontal line.

myininaya (myininaya):

L is the left intersection which is x=-2 U is the right intersection which is x=2 since f(2)=11

myininaya (myininaya):

\[\int\limits_{-2}^{2}(11-[x^3+3]) dx\] you should be able to do this part it always helps to draw a visual or see a visual for these types of problems

OpenStudy (anonymous):

integrate first than plug in upper/lower bounds?

OpenStudy (anonymous):

got 32 as a final answer.

OpenStudy (anonymous):

thanks for the help everybody.

OpenStudy (anonymous):

yeah frank but draw it first for it to make sense otherwise when you get into 3d volumes or surface area, it makes more sense

OpenStudy (anonymous):

it wont make as much sense*